# VoCoT: Unleashing Visually Grounded Multi-Step Reasoning in Large Multi-Modal Models

### v2

@@ B.1 Case Study
B.2 Limitation and Future Work
Although VoCoT can be applied to assist LMM in various tasks, there still exists limitations:
  • VoCalno is limited in terms of understanding certain visual-related concepts, such as spatial relationship and textures. We attribute this to insufficient training on vision-related data. In the future, richer data and text descriptions related to these concepts should be introduced in training LMMs.
  • Currently, VoCoT is not applicable to multi-image inputs. This issue could be addressed by including additional special tokens to enable VoCoT distinguish between objects in different images. We leave it as the future work.
  • Limited by the cost of data construction, VoCoT-Instruct can not be effectively scaled up currently. We hope to explore more efficient VoCoT data construction methods in the future.
  • VolCano is currently limited with respect to the model scale and input resolution. However, the experimental results reveal the potential of applying VoCoT in larger backbones, which is the future direction for the development of VolCano.

### v3

@@ B.8 Case Study
B.9 Discussion on Potential Social Impacts and Bias
We discuss potential issues and our solutions from the following perspectives:
  1. 1.
    Visual Bias: object categories in object detection datasets are unevenly distributed (mainly a long-tail distribution). To address the issue, we perform a balanced sampling of images based the included object categories. For the constructed dataset, we performed manual sampling and inspection and applied some filtering methods mentioned in Appendix A.1 to improve data quality. By checking the final constructed datasets, we did not observe any significant bias issues.
  2. 2.
    Misinformation: LMMs may produce erroneous information, namely hallucinations. The design of visually grounded representation in VoCoT aims to mitigate object hallucinations, and the experimental results validate the effectiveness.
  3. 3.
    Privacy issue: The images we adopt come from open-source and widely used datasets. Our construction method does not introduce additional privacy risks. Furthermore, we believe our object-centric method can be utilized to detect potential privacy issues in images. In future work, methods like RLHF will be used to guide VolCano to avoid detecting and analyzing data with potential privacy issues.
  4. 4.
    Beyond the above concerns. We utilize existing resources and there is no issue regarding potential personal information leakage and offensive content. The utilized tool, GPT-4V, also possess the capability to avoid generating offensive content. We manually checked the constructed dataset to ensure there is no such issues. We will continue to follow current responsible AI methods to monitor and alleviate our model and dataset for any biases or issues.
B.10 Discussion on the Use of Utilized and Presented Artifacts
In this work, we utilize existing artifacts including the data resources (GQA Hudson and Manning (2019), COCO Lin et al. (2014), and LVIS Gupta et al. (2019)), pre-trained models (CLIP Radford et al. (2021), Mistral Jiang et al. (2023), and Qwen2 Yang et al. (2024)), and existing datasets as listed in Table 1. All utilized artifacts are open-sourced to the research community. We carefully follow the license to use artifacts and ensure they are applicable for the research purpose. All utilized artifacts mainly focus on the English domain while Qwen2 and Mistral both possess multi-lingual capabilities. Please refer to the original resource for other information about the artifacts.
As for the artifacts we presented in this paper, including VoCoT-Instruct-80K and pre-trained VolCano and VolCanoQ2, we will release the data, code, and model weights to the community for research purpose. Our introduced artifacts are primarily designed for the English domain and will be extended to more languages. Our artifacts are designed with the principle of universality and fairness, without any preference for specific demographic groups.
B.11 Usage of AI Assistants
In this work, we mainly utilize GPT-4V as the AI assistants for preliminary exploration as in Figure 1, data transforming as in Section 2.2, and as a intelligent agent to judge the performance of models (Appendix B.4). Besides that, we utilize ChatGPT to help polish some parts of this paper.